I have 32 darts players they are in groups of 4 but playing as individuals,the groups of 4 needs to be different each week how do I work this out?
PS I am not that good at maths but the games due to start next week so any help would be great
Thanks for any help
Geoff

Answer by rothauserc(4718)

Number the dart players 1 through 32,
you are looking for the combination's of 32 players taken 4 at a time
this is 32! / (4! * (32-4)!) = (32*31*30*29) / (4*3*2*1) = 35960 possible combinations